Sprinkles285:

Almost 9 over 11 of the total electricity supplied to a town is used in farms and houses. Determine the decimal equivalent of 9 over 11 . Use words to explain if the decimal terminates or repeats. (10 points)

dukedadon:

The decimal is .81 repeating

dukedadon:

divide 9 and 11

Sprinkles285:

1.2?

kristen00001:

0.81 i think

dukedadon:

theres ya answer.. be sure to give medals (:

Sprinkles285:

umm... i don't think you're supposed to give me a direct answer.

Extrinix:

You aren't.

dukedadon:

here comes this guy lol.

Extrinix:

Okay so this is quite a simple question, all you have to do is divide 9 by 11, which the fraction is \(\dfrac{9}{11}\), and when divided gives you _____.

dukedadon:

i jus told him that .

Sprinkles285:

1.2?

Sprinkles285:

and the others are next: A newspaper reports the following changes in the temperature of a city over 4 days: 1.34 degrees Celsius, negative 5 over 7 degrees Celsius, −0.75 degrees Celsius, 4 over 9 degrees Celsius Write an expression to show the average daily change in the temperature of the city. Write the steps to solve the expression. Round your answer to the nearest hundredth. (10 points)

Extrinix:

So to find the average, let's start off by adding all degrees together and dividing by the amount of values, 4. \(\dfrac{1.34+ -\dfrac{5}{7}+-0.75+\dfrac{4}{9}}{4}\) = ? When added all together and divided, gives you roughly \(-8.50\).